OpenStudy (anonymous):

Can someone help me out with this problem please? Evaluate the following expression using the values given: Find 3x2 - y3 - y3 - z if x = 3, y = -2, and z = -5.

OpenStudy (opcode):

I take there are exponents? Such as: \[3x^2 - y^3 - y^3\]

OpenStudy (anonymous):

Yes thats correct.

OpenStudy (anonymous):

put the values of x y and z here

OpenStudy (opcode):

You are given the variables so plus them into the original equation. \(x = 3, y = -2, z = -5 \implies 3(3)^2 - (-2)^3 - (-2)^3 - (-5)\)

OpenStudy (opcode):

Do you know how to evaluate from here?: \[3(3)^2 - (-2)^3 - (-2)^3 - (-5)\] @Originality

OpenStudy (anonymous):

Sadly, I don't. :/

OpenStudy (opcode):

Okay, well do you understand everything I showed you before?

OpenStudy (anonymous):

\(3(3)^3-(-2)^3-(-2)^3-(-5)\) now what should you do as \((3)^3\) means multiply 3 to three times like 3*3*3 as it has powers 3=27 similarly, \((-2)^2\) has power 2 so multiply -2 two times like -2*-2=4

OpenStudy (opcode):

Yes, @ziqbal103 is correct, I assume he is taking lead now so I'll step back. :-)

OpenStudy (anonymous):

\oh bove is not 3^2 but 3^3 so it would be -2*-2*-2=-8

OpenStudy (anonymous):

and same way you take the next \(-(2)^3\)=-2*-2*-2=-8 and now it would be \(3(27)-(-8)-(-8)-(-5)\) keepit in your mind that always when two negative signs multiplied then it change into positive as done below \(81+8+8+5=102\)

OpenStudy (anonymous):

Is this for FLVS by any chance?